Usain Bolt can sprint 100 m in 10s. He accelerates
Answer #1

let the initial acceleration be a m/s^2.

initial speed=0 m/s


time of acceleration=3 seconds


final speed=initial speed+acceleration*time=0+a*3=3*a m/s

distance covered=initial speed*time+0.5*accceleration*time^2=0.5*a*9=4.5*a ma

for the next 7 seconds , he runs with uniform speed of 3*a m/s

then distance covered=speed*time=3*a*7=21*a m


total distance covered is 100 m

==>4.5*a+21*a=100

==>a=3.9216 m/s^2

part a:

his acceleration is 3.9216 m/s^2

part b:distance travelled in first 3 seconds=4.5*a=17.647 m

part c:

his final speed was 3*a=11.765 m/s